Fill in the missing numbers for the following income statement. (Input all amounts as positive values.)


  
  Sales $ 675,900
  Costs 430,800
  Depreciation 103,400


  EBIT $
  Taxes (30%)


  Net income $






Calculate the OCF.


  OCF $   


What is the depreciation tax shield?


  Depreciation tax shield $   
Answer #1

EBIT = sales - costs-depreciation = 675900 - 430800 - 103400 = $141,700


Net income = (1-tax rate ) EBIT = 0.7 * 141700 = $ 99,190


OCF = Net income - depreciation = 99,190 + 103400 = $202,590


Depreciation shield = tax rate * depreciation = 0.3 * 103400 = $ 31,020